# Quasi-Lindblad pseudomode theory for open quantum systems

### v1

@@ I Introduction
The conventional discrete mode representation is to directly discretize the continuous Hamiltonian [14, 15, 16] without introducing any dissipation. However, this discretization requires a large number of modes in the long-time limit, which scales linearly with the propagated time [14]. In most physical cases, the continuous bath provides dissipation to the system. It is thus natural to introduce an ansatz where the auxiliary bath dynamics is dissipative, in which case, the bath modes are referred to as pseudomodes [18, 19, 20, 26, 27, 28, 29]. The Liouvillian for the system and auxiliary bath can then be chosen to satisfy the constraint of physical dynamics, taking a Lindblad form and generating a completely positive trace-preserving (CPTP) dynamics. The corresponding BCF (assuming independent pseudomodes) is then a sum of complex exponential functions weighted by positive numbers. This BCF exactly represents the continuous spectral density as a positive weighted sum of Lorentzian functions. Unfortunately, many studies have found that this natural Lorentzian pseudomode expansion is inefficient, achieving only a slow rate of convergence as the number of pseudomodes increases [26, 25, 30]. Improving this convergence rate remains an active area of research [26, 30, 31, 32, 33, 34, 35].
@@
It is worth noting that the BCF alone does not uniquely determine the pseudomode formulation. Meanwhile, the system dynamics is uniquely determined by the BCF and is independent of the pseudomode formulation as long as the global dynamics can be simulated exactly (i.e., without further numerical approximations such as truncating the number of bosonic modes or even finite precision arithmetic operations). However, certain pseudomode formulations can be unstable if the corresponding Liouvillian has eigenvalues with positive real parts, leading to exponentially growing modes. This phenomenon has also been observed in other contexts, such as HEOM [47, 48, 49]. Therefore, in practice, the stability of the pseudomode formulation can indeed impact the quality of system dynamics. Nevertheless, with a careful choice of the pseudomode formulation that minimally violates the CP condition, we observe that the global dynamics can be stable across a range of physically relevant OQS in our simulations. We attribute this stability to the mixing of the divergent and dissipative parts of the global dynamics via the coherent terms in the Liouvillian.
@@ II.1 Problem setup: original unitary OQS
$C_{jj^{\prime}}^{B}(t,t^{\prime})=\operatorname{Tr}_{B}\left[\hat{B}_{j}(t)\hat{B}_{j^{\prime}}(t^{\prime})\hat{\rho}_{B}(0)\right]$ (2)
@@ II.4 Ansatz for quasi-Lindblad pseudomode
In addition to the enhanced flexibility, the use of complex exponential expressions also simplifies the numerical fitting of the BCF. In this work, we utilize the ESPRIT algorithm [37, 38], which has previously been reported as one of the most competitive methods for BCF fitting [44], to determine the complex exponents, $z_{k}$. We then perform a least-squares fitting of the complex weights $w_{k}$. We note that this simple unconstrained least-squares approach is possible because we do not restrict $w_{k}\geq 0$ as would be required in the conventional pseudomode approach. Once $w_{k}$ is determined, we choose the gauge for $V_{k}$ and $M_{k}$. In this limit, the gauge choice is parameterized as $V_{k}-iM_{k}=\kappa_{k}\sqrt{w_{k}},V_{k}+iM_{k}=\kappa_{k}^{*-1}\sqrt{w^{*}_{k}}$ where $\kappa_{k}\in\mathbb{C}$. Here, we follow a simple heuristic of minimizing the magnitude of $M_{k}$, as this term determines the extent of CP condition violation. We describe the result for the bosonic bath expression in Eq. 13 for $\bm{V}$ and $\bm{M}$, and its adaptation to the fermionic BCF expression is straightforward. The solution of $V_{k}$ and $M_{k}$ with minimum $|M_{k}|$ is then given by $V_{k}-iM_{k}=\sqrt{w_{k}}$ with $V_{k},M_{k}\in\mathbb{R}$, ($\kappa_{k}=1$) which we provide its proof in the SM [51]. We will examine other choices of $V_{k}$ and $M_{k}$ and their effect on the stability of the system-bath dynamics in Sec. IV. We note that [41] also provided a Lindblad-like pseudomode formulation derived from the HEOM with a similarity transformation. It is important to emphasize that this Lindblad-like equation represents one gauge choice for $\kappa_{k}$  [^3] The corresponding $\kappa_{k}$ for the Lindblad-like equation in [41] is $\kappa_{k}=\sqrt{\operatorname{Re}[w_{k}]/w_{k}}$. within our quasi-Lindblad pseudomode framework.
@@ V Conclusions
The pseudomode formulation provides an alternative framework to other numerical approaches to open quantum system dynamics, which have also used a complex-weighted sum of exponential functions to represent the BCF, such as the HEOM [40, 42, 41]. An important strength is that once the pseudomodes are determined, there exists a wide variety of numerical method that can be used to propagate the resulting master equation [55, 22, 23, 21, 31, 32, 68, 69].
@@
Acknowledgements.– This work is an equal collaboration between two SciDAC teams “Real-time dynamics of driven correlated electrons in quantum materials” and “Traversing the ‘death valley’ separating short and long times in non-equilibrium quantum dynamical simulations of real materials”, supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research and Office of Basic Energy Sciences, Scientific Discovery through Advanced Computing (SciDAC) program under Award Number DE–SC0022088 (G.P., C.Y., G.K.C.) and DE–SC0022198 (L.L., C.Y., Y.Z.). The work by Z.H. was supported by the Simons Targeted Grants in Mathematics and Physical Sciences on Moiré Materials Magic. G.K.C. and L.L. are Simons Investigators. We thank Zhiyan Ding, Joonho Lee, David Limmer, Vojtěch Vlček, Erika Ye, Lexing Ying for helpful discussions.
@@
$C_{j_{1},\cdots,j_{2n}}(t_{1},\cdots,t_{2n})=\sum_{\sigma\in S_{2n}}\prod_{i=1}^{n}C_{j_{\sigma(2i-1)},j_{\sigma(2i)}}(t_{\sigma(2i-1)},t_{\sigma(2i)}).$ (S6)
@@
$\hat{\widetilde{\rho}}_{S}(t)=\sum_{n=0}^{\infty}\frac{(-1)^{n}}{(2n)!}\int_{0}^{t}\cdots\int_{0}^{t}\sum_{j_{1}}\cdots\sum_{j_{n}}\sum_{\sigma\in S_{2n}}\mathcal{T}_{S}\left(\prod_{i=1}^{n}C_{j_{\sigma(2i-1)},j_{\sigma(2i)}}(t_{\sigma(2i-1)},t_{\sigma(2i)})\mathcal{S}_{j_{1}}(t_{1})\cdots\mathcal{S}_{j_{n}}(t_{n})\right)\hat{\rho}_{S}(0)dt_{1}\cdots dt_{n}$ (S7)
Thanks to the system time ordering superoperator, $\mathcal{T}_{S}$, $\mathcal{T}_{S}\left(\prod_{i=1}^{n}C_{j_{\sigma(2i-1)},j_{\sigma(2i)}}(t_{\sigma(2i-1)},t_{\sigma(2i)})\mathcal{S}_{j_{1}}(t_{1})\cdots\mathcal{S}_{j_{n}}(t_{n})\right)$ becomes the same operator for all $\sigma\in S_{2n}$. The number of elements of $S_{2n}$ is $(2n-1)!!$!, so Eq. S7 is reduced to,

### v2

@@ I Introduction
The conventional discrete mode representation is to directly discretize the continuous Hamiltonian [14, 15, 16] without introducing any dissipation. However, this discretization requires a large number of modes in the long-time limit, which scales linearly with the propagated time [14]. In most physical cases, the continuous bath provides dissipation to the system. It is thus natural to introduce an ansatz where the auxiliary bath dynamics is dissipative, in which case, the bath modes are referred to as pseudomodes [18, 19, 20, 26, 27, 28]. The Liouvillian for the system and auxiliary bath can then be chosen to satisfy the constraint of physical dynamics, taking a Lindblad form and generating a completely positive trace-preserving (CPTP) dynamics. The corresponding BCF (assuming independent pseudomodes) is then a sum of complex exponential functions weighted by positive numbers. This BCF exactly represents the continuous spectral density as a positive weighted sum of Lorentzian functions. Unfortunately, many studies have found that this natural Lorentzian pseudomode expansion is inefficient, achieving only a slow rate of convergence as the number of pseudomodes increases [26, 25, 29]. Improving this convergence rate remains an active area of research, with approaches including the introduction of additional intra-bath Hamiltonian terms [26, 29, 30, 31] or non-Hermitian pseudomode formulations [32, 33, 34, 35].
@@
It is worth noting that the BCF alone does not uniquely determine the pseudomode formulation. Meanwhile, the system dynamics is uniquely determined by the BCF and is independent of the pseudomode formulation as long as the global dynamics can be simulated exactly (i.e., without further numerical approximations such as truncating the number of bosonic modes or even finite precision arithmetic operations). However, certain pseudomode formulations can be unstable if the corresponding Liouvillian has eigenvalues with positive real parts, leading to exponentially growing modes. This phenomenon has also been observed in other contexts, such as HEOM [47, 48, 49, 50]. Therefore, in practice, the stability of the pseudomode formulation can indeed impact the quality of system dynamics. Nevertheless, with a careful choice of the pseudomode formulation that minimally violates the CP condition, we observe that the global dynamics can be stable across a range of physically relevant OQS in our simulations. We attribute this stability to the mixing of the divergent and dissipative parts of the global dynamics via the coherent terms in the Liouvillian.
@@ II.1 Problem setup: original unitary OQS
$C_{jj^{\prime}}^{B}(t,t^{\prime})=\operatorname{Tr}_{B}\left[\hat{B}_{j}(t)\hat{B}_{j^{\prime}}(t^{\prime})\hat{\rho}_{B}(0)\right],$ (2)
@@ II.4 Ansatz for quasi-Lindblad pseudomode
In addition to the enhanced flexibility, the use of complex exponential expressions also simplifies the numerical fitting of the BCF. In this work, we utilize the ESPRIT algorithm [37, 38], which has previously been reported as one of the most competitive methods for BCF fitting [44], to determine the complex exponents, $z_{k}$. We then perform a least-squares fitting of the complex weights $w_{k}$. We note that this simple unconstrained least-squares approach is possible because we do not restrict $w_{k}\geq 0$ as would be required in the conventional pseudomode approach. Once $w_{k}$ is determined, we choose the gauge for $V_{k}$ and $M_{k}$. In this limit, the gauge choice is parameterized as $V_{k}-iM_{k}=\kappa_{k}\sqrt{w_{k}},V_{k}+iM_{k}=\kappa_{k}^{*-1}\sqrt{w^{*}_{k}}$ where $\kappa_{k}\in\mathbb{C}$. Here, we follow a simple heuristic of minimizing the magnitude of $M_{k}$, as this term determines the extent of CP condition violation. We describe the result for the bosonic bath expression in Eq. 13 for $\bm{V}$ and $\bm{M}$, and its adaptation to the fermionic BCF expression is straightforward. The solution of $V_{k}$ and $M_{k}$ with minimum $|M_{k}|$ is then given by $V_{k}-iM_{k}=\sqrt{w_{k}}$ with $V_{k},M_{k}\in\mathbb{R}$, ($\kappa_{k}=1$) which we provide its proof in the SM [52]. We will examine other choices of $V_{k}$ and $M_{k}$ and their effect on the stability of the system-bath dynamics in Sec. IV.
There are several approaches to relax the CPTP condition in pseudomode representations. One approach involves a non-Hermitian pseudomode formulation [32, 33, 34, 35], where the system-bath coupling is given by a non-Hermitian Hamiltonian (without Lindblad coupling). This formulation results in real weights, $w_{k}\in\mathbb{R}$ [32, 35], or complex weights from a pair of pseudomodes [35]. Another approach, presented in [41], is a Lindblad-like pseudomode formulation derived from the HEOM with a similarity transformation. While this approach allows for complex weights, it is important to emphasize that the Lindblad-like equation in [41] represents one gauge choice for $\kappa_{k}$  [^3] The corresponding $\kappa_{k}$ for the Lindblad-like equation in [41] is $\kappa_{k}=\sqrt{\operatorname{Re}[w_{k}]/w_{k}}$. within our quasi-Lindblad pseudomode framework.
@@ V Conclusions
The pseudomode formulation provides an alternative framework to other numerical approaches to open quantum system dynamics, which have also used a complex-weighted sum of exponential functions to represent the BCF, such as the HEOM [40, 42, 41]. An important strength is that once the pseudomodes are determined, there exists a wide variety of numerical methods that can be used to propagate the resulting master equation [56, 22, 23, 21, 30, 31, 69, 70, 71].
@@
Acknowledgements.– This work is an equal collaboration between two SciDAC teams “Real-time dynamics of driven correlated electrons in quantum materials” and “Traversing the ‘death valley’ separating short and long times in non-equilibrium quantum dynamical simulations of real materials”, supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research and Office of Basic Energy Sciences, Scientific Discovery through Advanced Computing (SciDAC) program under Award Number DE–SC0022088 (G.P., C.Y., G.K.C.) and DE–SC0022198 (L.L., C.Y., Y.Z.). The work by Z.H. was supported by the Simons Targeted Grants in Mathematics and Physical Sciences on Moiré Materials Magic. G.K.C. and L.L. are Simons Investigators. We thank Zhiyan Ding, Joonho Lee, David Limmer, Vojtěch Vlček, Erika Ye, Lexing Ying, and Neill Lambert for helpful discussions.
@@
Note added.– During submission of this manuscript, a related paper [75] appeared, which discusses analytical numerical scaling and provides different numerical fitting schemes.
@@
$C_{j_{1},\cdots,j_{2n}}(t_{1},\cdots,t_{2n})=\sum_{\sigma\in\Pi_{2n}}\prod_{i=1}^{n}C_{j_{\sigma(2i-1)},j_{\sigma(2i)}}(t_{\sigma(2i-1)},t_{\sigma(2i)}).$ (S6)
@@
$\hat{\widetilde{\rho}}_{S}(t)=\sum_{n=0}^{\infty}\frac{(-1)^{n}}{(2n)!}\int_{0}^{t}\cdots\int_{0}^{t}\sum_{j_{1}}\cdots\sum_{j_{n}}\sum_{\sigma\in\Pi_{2n}}\mathcal{T}_{S}\left(\prod_{i=1}^{n}C_{j_{\sigma(2i-1)},j_{\sigma(2i)}}(t_{\sigma(2i-1)},t_{\sigma(2i)})\mathcal{S}_{j_{1}}(t_{1})\cdots\mathcal{S}_{j_{n}}(t_{n})\right)\hat{\rho}_{S}(0)dt_{1}\cdots dt_{n}$ (S7)
Thanks to the system time ordering superoperator, $\mathcal{T}_{S}$, $\mathcal{T}_{S}\left(\prod_{i=1}^{n}C_{j_{\sigma(2i-1)},j_{\sigma(2i)}}(t_{\sigma(2i-1)},t_{\sigma(2i)})\mathcal{S}_{j_{1}}(t_{1})\cdots\mathcal{S}_{j_{n}}(t_{n})\right)$ becomes the same operator for all $\sigma\in\Pi_{2n}$. The number of elements of $\Pi_{2n}$ is $(2n)!!$!, so Eq. S7 is reduced to,